# The Quantum Measurement Problem: A Review of Recent Trends

### v2

@@ 5.1 Many-Worlds Interpretation
The lack of information on the outcomes is not a problem for some interpretations. For instance, the Many-Worlds Interpretation (or Relative State Interpretation or, still, Everettian Quantum Mechanics) proposes that the remaining outcomes after decoherence correspond to independent universe branches [45]. Observers existing in each branch perceive different outcomes because their quantum state, entangled with the molecular quantum state, also branches. In one branch, the combined molecule-observer quantum state may be and in another, . No branches with superpositions of both states exist because of decoherence toward the environment (surrounding the system and the observer) suppressing them [46]. Thus, the Many-Worlds Interpretation posits the existence of a single quantum universe composed of non-communicating classical branches.
@@
For the Many-Worlds Interpretation, the quantum state is a complete description of reality (no hidden variables are required), exclusively evolving according to the Schrödinger Equation (collapse is only an illusion caused by the universe branching) [47, 48]. Thus, the Many-Worlds Interpretation is consistent with standard quantum mechanics and requires no theory modification. However, it remains an interpretation rather than a theory until it gains experimental testability. A significant challenge is the technical infeasibility of maintaining and measuring a quantum system in a coherent superposition of orthogonal states, especially for large systems. This difficulty explains why superpositions of macroscopic states cannot be observed directly due to rapid decoherence. Susskind and Aaronson demonstrated this limitation using methods from quantum computational complexity, naming it the quantum-hard problem [49, 50].
@@
A puzzling feature of the Many-Worlds interpretation is understanding how the deterministic branching relates to the wave function amplitudes and the Born rule. Carroll and Sebens [51] propose that this relation is established after decoherence occurs but before the measurement is registered. In this period, if the observer wants to estimate which branch they are, their rational credence directly leads to the Born rule. Thus, the probabilistic character of quantum mechanics would be a subjective aspect arising from the observer’s self-locating uncertainty. Carroll and Sebens’ derivation of the Born rule is only strictly valid if the ratios of branch amplitudes are the square roots of rational numbers. However, they argue that the large number of branches in practical situations always allows this condition to be approximately satisfied. This approximation also allows for experimental tests of their hypothesis by searching for small deviations in the Born rule. Kent [52] forcefully criticized Carroll and Sebens’ argument regarding their conceptual foundation and mathematical rigor. His main criticism is that an underlying ontology supporting Carroll-Sebens’ model would require a unique and objective branching scheme.
@@
The Many-Worlds Interpretation may seem exotic to non-specialists due to its exuberant ontology. Despite this perception, it is a serious contender among the proposed solutions to the Measurement Problem. It is a popular interpretation among physicists, with prominent supporters like Schlosshauer [24], Carroll [47, 48, 51], Zurek [46], Zeh [42], and Deutsch promoting different variants. However, it also faces outstanding critics, like Penrose [54] and Albert. For Landsmann, the Many-Worlds Interpretation would be “acceptable only if truly everything else has failed” [39, Ch 11].
@@ 5.2 Epistemic Interpretations
Epistemic interpretations claim that a quantum state is no more than a mathematical tool to predict outcomes. This is the stance of interpretations such as the Relational Interpretation [55, 56, 57], Quantum Information-Theoretic approaches [58], Quantum Bayesianism [59], and Indivisible Stochastic Process Theory [60], which are discussed next.
@@ 5.2.1 Relational Interpretation
Rovelli draws attention to the fact that this state dependence on a specific observer is not new in physics. It is analogous, for instance, to a velocity dependence on the particular reference frame where it is measured. The novelty in quantum mechanics is that the non-commutative nature of quantum variables implies that not all properties can be sharply defined simultaneously [57], limiting the completeness of information derivable from a single interaction.
@@
This perspective may offer new insights into the Measurement problem, particularly in Wigner’s friend scenarios, where the possibility of conflicting accounts among observers becomes natural if each observer constitutes its own QFR [61, 62]. Section 5.3.2 discusses how Giacomini and Brukner use QFRs to challenge Pennrose’s conjecture on gravity-induced collapse [63].
@@ 5.2.2 Information-Theoretic and QBism Views
Some no-go theorems are sometimes cited as challenging epistemic views, particularly the Pusey-Barrett-Rudolph (PBR) theorem [66]. The theorem shows that, under certain assumptions, quantum states cannot be interpreted merely as statistical knowledge about underlying physical states—that is, hidden variables (see Section 5.4). Instead, the quantum state must be part of the physical reality. However, this conclusion applies only to models that posit such hidden variables. Interpretations like QBism, which deny the existence of underlying ontic states altogether, fall outside the scope of the theorem [67, 68]. Rovelli has made similar clarifications for relational interpretations as well [69].
@@
A common misunderstanding is to conflate QBism with older subjectivist interpretations, like that of London and Bauer [73], who suggested that collapse results from an observer becoming aware of an outcome. However, QBism takes a different stance: the measurement outcome is not something the agent discovers; it is the agent’s experience. No underlying real wave function remains untouched—it is a theory about actions and consequences, not objective ontologies [74].
@@ 5.2.3 Indivisible Stochastic Process Theory
Barandes has recently proposed the Indivisible Stochastic Processes theory, where classical configurations of particles, fields, or qubits evolve via non-Markovian dynamics [60, 75]. In this type of dynamics, which recovers all results from conventional quantum mechanics, transition probabilities cannot be estimated at intermediate times. In Barandes’ approach, wave functions are considered predictive tools only, and the Hilbert space of quantum mechanics is demoted to a convenient instrumentalist framework without an ontological basis.
@@ 5.3.1 Agnostic Objective Collapse
Opposed to Many-Worlds and Epistemic Interpretations, Objective Collapse Theories consider that collapse is a physical process, and that standard quantum mechanics is still incomplete [5, 8]. Thus, these theories attempt to change the Schrödinger Equation to account for the collapse. This subsection deals with agnostic objective collapse, that is, theories that don’t attribute a cause to the collapse. In the following subsection, we will look at gravity-induced collapse.
@@
In addition to the usual Hamiltonian term, the right-hand side contains two additional terms that tend to localize the state in one of the eigenstates of the operators , as illustrated in Figure 5 for . The collapse is stochastic and controlled by a Wiener process (with and ), which recovers the Born rule. A positive real constant tells how wide the noise is and how fast the collapse is. The equation is nonlinear due to the terms. However, the evolution of the ensemble density is linear thanks to the quadratic last term [81]. This linearity is required to avoid superluminal information exchange.
Different objective collapse theories have been proposed, corresponding to specific choices of operators , including the Hamiltonian operator [82], position operators (GRW after the initials of the authors [83]) density number operators (CSL for Continuous Spontaneous Localization [77]), and mass density operators [84] (which is discussed in Section 5.3.2). These specific models are thoroughly reviewed in Ref. [5, 9].
@@ 5.3.2 Gravity-Induced Objective Collapse
The Diósi-Penrose model has attracted criticism on multiple fronts. Gao [87] raised several issues on the Penrose conjecture. His main point is that Penrose’s argument by analogy with the conventional quantum mechanical time-uncertainty principle is not rigorous enough. Moreover, he is also right to point out that although we commonly refer to the Di’osi-Penrose model, Diósi and Penrose theories are fundamentally distinct, arriving at the same collapse time through independent, though conceptually related, arguments. Diósi himself [88] notes their similarities and differences and suggests that further work is needed to clarify whether the two frameworks are compatible.
Giacomini and Bruckner [63] also challenged Penrose’s conjecture that quantum superposition conflicts with Einstein’s equivalence principle. Using the Quantum Reference Frames (QRF) formalism (see also 5.2.1), they showed that it’s possible to generalize the principle (or, more precisely, always to make the metric field locally Minkowskian) even in the presence of quantum superpositions of massive bodies.
Additionally, we performed atomistic simulations to show that the Diósi-Penrose collapse time reaches a minimum value independent of superposition displacement and, in some instances, can be surprisingly long [1]. Figurato et al. [89] has also independently concluded that the model could have too long collapse times for macroscopic bodies. Beyond these theoretical concerns, experimental tests have also challenged the model. Donadi et al. [90] and Arnquist et al. [91] conducted underground experiments to test gravity-related collapse mechanisms. Their results placed stringent constraints on the model’s parameters (see Section 5.3.3).
@@
Oppenheim has recently proposed another hypothesis that attributes wave function collapse to the influence of gravity. Indeed, his Postquantum Theory of Classical Gravity does not focus on the measurement problem and has a more ambitious goal of reconciling gravity and quantum mechanics [93]. Oppenheim’s central hypothesis is that gravity is a classical field. To ensure classical gravity can interact coherently with quantum matter fields, Oppenheim developed a Classical Quantum (CQ) Dynamics framework, which circumvents the usual obstacles encountered by semiclassical models, such as violation of the superposition principle due to nonlinearities in the equations of motion [93, 94, 95].
@@
Even though stochastic trajectories are introduced as a means to propagate the density evolution (5.3.2), Oppenheim and co-authors assign them ontological significance [97]. A stochastic trajectory represents the system collapsing to a specific state, just as it is also done in the other objective collapse theories discussed above. Oppenheim’s theory implies that collapse does not require measurement, as it occurs naturally due to interactions with a classical field [97].
@@ 5.3.3 Experimental Assessment of Objective Collapse Theories
These efforts signal a shift in the status of objective collapse theories from speculative constructs to empirically testable frameworks. Among them, CSL stands out for its tunable parameters and broad compatibility with experimental platforms, while the Diòsi-Penrose and Oppenheim models offer more rigid, gravity-motivated alternatives. For example, Angeli et al. recently demonstrated that if gravity is both classical and local, as in Oppenheim’s proposal (see Section 5.3.2), then it must induce diffusion in quantum systems, a phenomenon potentially observable with large-scale mechanical probes such as torsion pendulums [102].
Altogether, these developments highlight the growing entanglement of collapse models with experimental physics. Whether or not they ultimately capture physical reality, they provide concrete benchmarks for testing the limits of quantum theory and may yet point toward deeper structures beyond it.
@@ 5.4.1 Physical Restrictions on Hidden Variables
Hidden variable theories (also known as ontic theories) explore the hypothesis that the quantum state may not be a complete description of the system and that a set of additional (hidden) variables may contain further information (accessible or not) [103]. The hidden variables discussion is where the Measurement Problem spills the most on other fundamental questions of quantum mechanics, like nonlocality in entangled systems, the completeness of quantum mechanics, and the existence of properties independently of measurements.
The hidden variable theories are usually framed within the limits of two no-go theorems and a couple of physically motivated restrictions. The theorems are the Bell and the Kochen-Specker ones. Bell’s theorem states that no local hidden-variable theory can be entirely compatible with quantum mechanics [104, Ch 2] [105]. Local in this context means that “the result of a measurement on one system be unaffected by operations on a distant system with which it has interacted in the past” [104, Ch 2].
@@
The situation seems simple from there. For instance, the validity of Bell’s Theorem—which experiments have vastly supported [108]—seems to imply that physical theories must either be non-local or not have properties defined at all times. However, no-go theorems are as strong as their underlying hypotheses, often deeply hidden in the deduction. The restrictions imposed by these theorems and physical principles have been constantly challenged [109, 110] (except for the non-signaling principle, which is likely the only consensual rule in the game). Relaxing statistical independence (sometimes loosely called the free-will principle) in Bell’s theorem, for example, may allow for building consistent local and realist theories. We will return to this point in Section 5.5.
@@ 5.4.2 Bohmian Mechanics
Most of the practical development in hidden-variable theories happens in the framework of Bohmian Mechanics, which posits the existence of corpuscules moving in a non-Newtonian way and guided by a wave function [111]. The corpuscule’s velocities are determined by a velocity field (the de Broglie-Bohm Equation) dependent on the wave function, which, in turn, follows the Schrödinger evolution. This deterministic time evolution in Bohmian Mechanics automatically solves the problem of outcomes. Although Bohmian Mechanics’ experimental predictions are equivalent to quantum mechanics [112], it is not discarded that they could be experimentally tested [113].
Poirier has proposed an alternative theory that prescribes non-Newtonian trajectories but entirely discards wave functions [114, 115]. In this case, the quantum system is described by an ensemble of interacting deterministic real-valued trajectories with well-defined positions and momenta. This type of theory inscribes itself into the tradition of hydrodynamic interpretation of quantum mechanics, inaugurated by Madelung right after Schrödinger’s original work [116][117, p. 222]. In Madelung’s interpretation, quantum mechanics is modeled as a quantum fluid [118], which can be understood as many interacting classical worlds [119].

### v3

@@ 5.1 Many-Worlds Interpretation
The lack of information on the outcomes is not a problem for some interpretations. For instance, the Many-Worlds Interpretation (or Relative State Interpretation or, still, Everettian Quantum Mechanics) proposes that the remaining outcomes after decoherence correspond to independent universe branches [49]. Observers existing in each branch perceive different outcomes because their quantum state, entangled with the molecular quantum state, also branches. In one branch, the combined molecule-observer quantum state may be and in another, . No observed branches with superpositions of both states exist because of decoherence toward the environment (surrounding the system and the observer) suppressing them [50]. Thus, the Many-Worlds Interpretation posits the existence of a single quantum universe composed of non-communicating classical branches.
@@
For the Many-Worlds Interpretation, the quantum state is a complete description of reality (no hidden variables are required), exclusively evolving according to the Schrödinger equation (collapse is only an illusion caused by the universe branching) [51, 52]. Thus, the Many-Worlds Interpretation is consistent with standard quantum mechanics and requires no modification to the theory. However, it remains an interpretation rather than a testable theory, since it currently makes no unique experimental predictions. A significant challenge is the technical infeasibility of maintaining and measuring a quantum system in a coherent superposition of orthogonal states, especially for large systems. This difficulty explains why superpositions of macroscopic states cannot be observed directly due to rapid decoherence. Susskind and Aaronson demonstrated this limitation using methods from quantum computational complexity, naming it the quantum-hard problem [53, 54].
@@
A puzzling feature of the Many-Worlds Interpretation is understanding how the deterministic branching relates to the wave function amplitudes and the Born rule. Carroll and Sebens [55] propose that this relation is established after decoherence occurs but before the measurement is registered. In this period, if the observer wants to estimate which branch they are, their rational credence directly leads to the Born rule. Thus, the probabilistic character of quantum mechanics would be a subjective aspect arising from the observer’s self-locating uncertainty. Carroll and Sebens’ derivation of the Born rule is only strictly valid if the ratios of branch amplitudes are the square roots of rational numbers. However, they argue that the large number of branches in practical situations always allows this condition to be approximately satisfied. This approximation also enables experimental tests of their hypothesis by searching for minor deviations from the Born rule. Kent [56] sharply questioned the conceptual and mathematical soundness of Carroll and Sebens’ account, arguing that their notion of self-locating uncertainty presupposes an objective and well-defined branching of the wave function—something not provided by the Everettian framework itself.
@@
The Many-Worlds Interpretation may seem exotic to non-specialists due to its exuberant ontology. Despite this perception, it is a serious contender among the proposed solutions to the Measurement Problem. It is a popular interpretation among physicists, with prominent supporters like Carroll [51, 52, 55], Tegmark [58], Zeh [59, 46, 60], and Deutsch—though their formulations vary in how explicitly Everettian they are and in the extent of their commitment to the metaphysics of many worlds. However, it also faces outstanding critics, like Penrose [61] and Albert. For Landsmann, the Many-Worlds Interpretation would be “acceptable only if truly everything else has failed” [43, Ch 11].
@@ 5.2 Epistemic Interpretations
Epistemic interpretations claim that a quantum state is no more than a mathematical tool to predict outcomes. This stance is shared by interpretations such as the Relational Interpretation [62, 63, 64], Quantum Information-Theoretic approaches [65], Quantum Bayesianism [66], and Indivisible Stochastic Process Theory [67], which are discussed next.
@@ 5.2.1 Relational Interpretation
Rovelli draws attention to the fact that this state dependence on a specific observer is not new in physics. It is analogous, for instance, to the velocity dependence on the particular reference frame in which it is measured. The novelty in quantum mechanics is that the non-commutative nature of quantum variables implies that not all properties can be sharply defined simultaneously [64], limiting the completeness of information derivable from a single interaction.
@@
This perspective may offer new insights into the Measurement Problem, particularly in Wigner’s friend scenarios, where the possibility of conflicting accounts among observers becomes natural if each observer constitutes its own QRF [68, 69]. Section 5.3.2 discusses how Giacomini and Brukner use QRFs to challenge Penrose’s conjecture on gravity-induced collapse [70].
@@ 5.2.2 Information-Theoretic and QBism Views
Some no-go theorems are sometimes cited as challenging epistemic views, particularly the Pusey-Barrett-Rudolph (PBR) theorem [73]. It shows that, under certain assumptions, quantum states cannot be interpreted merely as statistical knowledge about underlying physical states—that is, hidden variables (see Section 5.4). Instead, the quantum state must be part of physical reality. However, this conclusion applies only to models that posit such hidden variables. Interpretations like QBism, which deny the existence of underlying ontic states altogether, fall outside the scope of the theorem [74, 75]. Rovelli has made similar clarifications for relational interpretations as well [76].
@@
A common misunderstanding is to conflate QBism with older subjectivist interpretations, such as that of London and Bauer [80], who suggested that collapse results from an observer becoming aware of an outcome. However, QBism takes a different stance: the measurement outcome is not something the agent discovers; it is the agent’s experience. No underlying real wave function remains untouched—it is a theory about actions and consequences, not objective ontologies [81].
@@ 5.2.3 Indivisible Stochastic Process Theory
Barandes has recently proposed the Indivisible Stochastic Process Theory, where classical configurations of particles, fields, or qubits evolve via non-Markovian dynamics [67, 82]. In this type of dynamics, which recovers all results from conventional quantum mechanics, transition probabilities cannot be estimated at intermediate times. In Barandes’ approach, wave functions are considered predictive tools only, and the Hilbert space of quantum mechanics is demoted to a convenient instrumentalist framework without an ontological basis.
@@ 5.3.1 Agnostic Objective Collapse
Opposed to Many-Worlds Interpretation and epistemic interpretations, objective collapse theories consider that collapse is a physical process, and that standard quantum mechanics is still incomplete [5, 8]. Thus, these theories attempt to change the Schrödinger equation to account for the collapse. This subsection deals with agnostic objective collapse, that is, theories that do not attribute a cause to the collapse. In the following subsection, we will look at gravity-induced collapse.
@@
In addition to the usual Hamiltonian term, the right-hand side contains two additional terms that tend to localize the state in one of the eigenstates of the operators , as illustrated in Figure 5 for . The collapse is stochastic and controlled by a Wiener process (with and ), which, for the ensemble density, the statistical distribution of outcomes follows the Born rule. A positive real constant sets the noise strength (diffusion rate), hence the collapse timescale. The equation is nonlinear due to the terms. However, the evolution of the ensemble density is linear thanks to the quadratic last term [88]. This linearity is required to avoid superluminal information exchange.
Different objective collapse theories have been proposed, corresponding to specific choices of operators , including the Hamiltonian operator [89], position operators (GRW after the initials of the authors [90]), density number operators (CSL for Continuous Spontaneous Localization [84]), and mass density operators [91] (which is discussed in Section 5.3.2). These specific models are thoroughly reviewed in Ref. [5, 9].
@@ 5.3.2 Gravity-Induced Objective Collapse
The Diósi-Penrose model has attracted criticism on multiple fronts. Gao [94] raised several issues on the Penrose conjecture. His main point is that Penrose’s argument by analogy with the conventional quantum mechanical time-uncertainty principle is not rigorous enough. Moreover, he is also right to point out that although we commonly refer to the Diósi-Penrose model, Diósi and Penrose theories are fundamentally distinct, arriving at the same collapse time through independent, though conceptually related, arguments. Diósi himself [95] notes their similarities and differences and suggests that further work is needed to clarify whether the two frameworks are compatible.
Giacomini and Bruckner [70] also challenged Penrose’s conjecture that quantum superposition conflicts with Einstein’s equivalence principle. Using the Quantum Reference Frames (QRF) formalism (see also 5.2.1), they showed that it is possible to generalize the principle (or, more precisely, always to make the metric field locally Minkowskian) even in the presence of quantum superpositions of massive bodies.
Additionally, our atomistic simulations showed that the Diósi-Penrose collapse time reaches a minimum value independent of superposition displacement and, in some instances, can be surprisingly long [1]. Figurato et al. [96] have also independently concluded that the model could have too long collapse times for macroscopic bodies. Beyond these theoretical concerns, experimental tests have also challenged the model. Donadi et al. [97] and Arnquist et al. [98] conducted underground experiments to test gravity-related collapse mechanisms. Their results placed stringent constraints on the model’s parameters (see Section 5.3.3).
@@
Oppenheim has recently proposed another hypothesis that attributes wave function collapse to the influence of gravity. Indeed, his Postquantum Theory of Classical Gravity does not focus on the Measurement Problem and has a more ambitious goal of reconciling gravity and quantum mechanics [100]. Oppenheim’s central hypothesis is that gravity is a classical field. To ensure classical gravity can interact coherently with quantum matter fields, Oppenheim and colleagues developed a Classical Quantum (CQ) Dynamics framework, which circumvents the usual obstacles encountered by semiclassical models, such as violation of the superposition principle due to nonlinearities in the equations of motion [100, 101, 102].
@@
Even though stochastic trajectories are introduced as a means to propagate the density evolution (5.3.2), Oppenheim and co-authors assign them ontological significance [104]. A stochastic trajectory represents the system collapsing to a specific state, just as it is also done in the other objective collapse theories discussed above. Their theory implies that collapse does not require measurement, as it occurs naturally due to interactions with a classical field [104].
@@ 5.3.3 Experimental Assessment of Objective Collapse Theories
These efforts signal a shift in the status of objective collapse theories from speculative constructs to empirically testable frameworks. Among them, CSL stands out for its tunable parameters and broad compatibility with experimental platforms, while the Diósi-Penrose and Oppenheim models offer more rigid, gravity-motivated alternatives. For example, Angeli et al. recently demonstrated that if gravity is both classical and local, as in Oppenheim’s proposal (see Section 5.3.2), then it must induce diffusion in quantum systems, a phenomenon potentially observable with large-scale mechanical probes such as torsion pendulums [109].
Altogether, these developments underscore the increasing interconnection of collapse models with experimental physics. Whether or not they ultimately capture physical reality, they provide concrete benchmarks for testing the limits of quantum theory and may yet point toward deeper structures beyond it.
Despite the increasing experimental scrutiny of objective-collapse models, the idea that the Schrödinger equation may provide only an incomplete description of physical evolution continues to inspire new proposals. A recent one by Dick reframes the Measurement Problem as an interaction problem [110]. In this view, every inelastic energy exchange between systems acts as a mutual observation, interrupting the smooth Schrödinger evolution and producing definite outcomes. The wave function is interpreted as an epistemic superposition of possible ontic outcomes—elastic channels describing continuous evolution, inelastic ones involving discontinuous quantum jumps. Although this approach introduces no explicit modification of the Schrödinger equation, it assumes that the equation captures only the evolution of relative probability amplitudes, not the actual transitions that occur during energy exchange. In this sense, Dick’s proposal can be regarded as a minimalist form of objective collapse, locating the breakdown of unitarity within ordinary interactions rather than in new stochastic dynamics.
@@ 5.4.1 Physical Restrictions on Hidden Variables
Hidden variable theories (also known as ontic theories) explore the hypothesis that the quantum state may not be a complete description of the system and that a set of additional (hidden) variables may contain further information (accessible or not) [111]. The hidden variables discussion is where the Measurement Problem spills the most on other fundamental questions of quantum mechanics, such as nonlocality in entangled systems, the completeness of quantum mechanics, and the existence of properties independently of measurement.
Hidden variable theories are typically framed within the limits of two no-go theorems and a few physically motivated restrictions. The theorems are the Bell and the Kochen-Specker ones. Bell’s theorem states that no local hidden-variable theory can be entirely compatible with quantum mechanics [112, Ch 2] [113]. Local in this context means that “the result of a measurement on one system be unaffected by operations on a distant system with which it has interacted in the past” [112, Ch 2].
@@
The situation seems simple from there. For instance, the experimental violations of Bell’s inequalities [116]—confirming the predictions of quantum mechanics—imply that physical theories must either be non-local or not have properties defined at all times. However, no-go theorems are as strong as their underlying hypotheses, often deeply hidden in the deduction. The restrictions imposed by these theorems and physical principles have been constantly challenged [117, 118] (except for the non-signaling principle, which is likely the only consensual rule in the game). Relaxing statistical independence (sometimes loosely called the free-will principle) in Bell’s theorem, for example, may allow for building consistent local and realist theories. We will return to this point in Section 5.5.
@@ 5.4.2 Bohmian Mechanics
Most of the practical development in hidden-variable theories happens in the framework of Bohmian Mechanics, which posits the existence of corpuscles moving in a non-Newtonian way and guided by a wave function [119]. The corpuscle’s velocities are determined by a velocity field (the de Broglie-Bohm Equation) dependent on the wave function, which, in turn, follows the Schrödinger evolution. This deterministic time evolution in Bohmian Mechanics automatically solves the problem of outcomes. Although Bohmian Mechanics’ experimental predictions are equivalent to quantum mechanics [120], it is not discarded that they could be experimentally tested [121]. A striking example is the experiment by Sharoglazova et al. [122], who investigated the energy-speed relationship of wave-guided photons in a microcavity. They observed that for evanescent states within a potential step, lower-energy particles exhibited higher speeds. This result is in tension with standard Bohmian Mechanics, which predicts zero speed for such particles.
Poirier has proposed an alternative theory that prescribes non-Newtonian trajectories but entirely discards wave functions [123, 124]. In this case, the quantum system is described by an ensemble of interacting deterministic real-valued trajectories with well-defined positions and momenta. This type of theory inscribes itself into the tradition of hydrodynamic interpretation of quantum mechanics, inaugurated by Madelung right after Schrödinger’s original work [125][126, p. 222]. In Madelung’s interpretation, quantum mechanics is modeled as a quantum fluid [127], which can be understood as many interacting classical worlds [128].